OpenStudy (anonymous):

Let f(x) = ln(9 − x). (a) Find the domain of f. −infinity < x < infinity x ≤ 9 x ≥ 9 x < 9 x > 0 (b) Find the range of f. y > 0 y > 4 −infinity < y < infinity y ≥ 4 y < 9 (b) Find f^−1(x). f^−1(x) =

OpenStudy (anonymous):

for a it should be x>0

OpenStudy (anonymous):

wait i mean x<=9

ganeshie8 (ganeshie8):

ln(0) is also undefined

OpenStudy (anonymous):

huh what happened to his answer??

OpenStudy (zzr0ck3r):

ln(x) domain is all x > 0 so ln(9-x) domain is when 9-x> 0 -x> -9 x<9

OpenStudy (zzr0ck3r):

y = ln(9-x) so e^y = 9-x so e^y-9 = -x x = 9-e^y so the range is all real numbers

OpenStudy (zzr0ck3r):

and the inverse I already showed but rename x to y and y to x y = 9-e^x

OpenStudy (anonymous):

can u explain this to me, my teacher is horrible :( y = ln(9-x) so e^y = 9-x so e^y-9 = -x x = 9-e^y

OpenStudy (zzr0ck3r):

The way I find the range of something is first take the inverse, then I see what are the domain restrictions on the inverse. this tells me the range.

OpenStudy (zzr0ck3r):

so like y = 1/x has a domain of all x such that x != 0 and the range... y = 1/x so x = 1/y now the range is all the values of y, and we see clearly that y cant be 0, so the range is all numbers x such that x != 0

OpenStudy (anonymous):

where do u get the e^y tho??

OpenStudy (zzr0ck3r):

y = ln(9-x) e^y = e^ln(9-x) e^y = 9-x make sense?

OpenStudy (zzr0ck3r):

if y = lnx then x = e^y by definition of ln(x)